# Rethinking Fine-Tuning when Scaling Test-Time Compute: Limiting Confidence Improves Mathematical Reasoning

### v3

@@ 1 Introduction
  • We identify a misalignment between standard fine-tuning with CE loss and the pass@N coverage metric at test time (Section 4.1).
  • We develop and experimentally verify a framework that suggests this misalignment arises from overconfidence induced by training on CE loss (Sections 4.2 and 4.3).
  • We propose a loss function that directly optimizes the pass@N coverage metric, demonstrating consistent improvement over the CE loss objective and achieving superior accuracy frontiers on MATH and MiniF2F (Sections 5.1, 5.2 and 5.3).
  • We extend our algorithm to more complex test-time scenarios including searching over proof-trees of varying shapes to improve automated theorem proving, and answering math questions using Chain-of-Thought reasoning traces, demonstrating improved mathematical reasoning in both cases (Sections 5.3 and 5.4).
@@ 2 Related Works
Test-time compute and pass@N strategy. A growing body of work explores how training and test-time strategies interact to shape model performance. Jones [15] highlight a tradeoff between train and test-time compute in a controlled board game setting. Brown et al. [16] identify an exponentiated power law between coverage and number of in-context samples, while Snell et al. [1] explore compute-optimal strategies for scaling test-time inference. Our paper focuses primarily on the pass@N test-time strategy. Gui et al. [17] show that, in the best-of-N setting, the sample distribution is well-aligned with the reward model, and alignment methods have been proposed to distill this distribution and reduce sampling costs [17, 18]. Closely related to our work, Chow et al. [14] derives a similar training objective for reinforcement learning (RL) to directly optimize for the best-of-N test-time strategy. Li et al. [19] argue that fine-tuning with cross-entropy loss limits output diversity and propose a maximum entropy-based method to address this issue. Our work similarly identifies limitations of cross-entropy when scaling test-time compute, but focuses on the alignment of training and test-time strategies. In addition, Yue et al. [20] argue that fine-tuning with Group Relative Policy Optimization (GRPO) does not improve reasoning capabilities beyond the base model when evaluated with pass@N strategy. We will show that GRPO training may also lead to overconfidence in Figure 3.
@@ 3 Problem setup
Minimizing the CE loss in Equation 1 is equivalent to maximizing the pass@$1$ metric $\mathcal{C}_{\mathcal{D}}^{1}$ on a training set $\mathcal{D}$. But if we scale up test-time compute so that the pass@N metric $\mathcal{C}_{\mathcal{D}}^{N}$ on a test set $\mathcal{D}$ for $N\gg 1$ is the relevant performance metric, is $\mathcal{L}_{\mathrm{CE}}$ still a good training loss, or can we do better?
@@ 4.1 The CE loss induces overfitting for pass@N
To understand the impact of training with CE loss on pass@N test performance, we fine-tune Llama-3-8B-base [40] on the MATH [41] dataset. We start from the base model rather than Llama-3-8B-Instruct to avoid potential leakage of the MATH dataset into Llama-3-8B-Instruct through post-training. We follow Lightman et al. [22] and use $12,000$ problems for training and the remaining $500$ for testing. Here we train the model to provide a direct answer without a reasoning trace. We will discuss training with CoT in Section 5.4. We confirm that the results presented in this section also hold for a larger model, Llama-3-70B-base, and direct readers to Section C.3 for additional results.
@@
Table 1 reveals that the pass@N performance $\mathcal{C}_{\mathcal{D}}^{N}$ on a test set monotonically increases with the number of training epochs only for $N=1$, when minimizing CE loss is equivalent to maximizing $\mathcal{C}_{\mathcal{D}}^{1}$. However, for $N\geq 16$, minimizing CE loss during training does not monotonically increase $\mathcal{C}_{\mathcal{D}}^{N}$ at test; indeed for $N\geq 256$, pass@N test performance, remarkably, monotonically decreases with the number of training epochs, despite the fact that pass@1 performance monotonically increases. This effectively corresponds to a novel type of overfitting, in which test performance degrades over training, likely due to a mismatch between the test time (pass@N) and training time (pass@1) strategies.
@@ 4.2 Overfitting, confidence, and explore-exploit tradeoff
However, Lemma 4.1 can fail when there is more than one problem. To understand this in a simple setting, consider a test set with two problems $x_{1}$ and $x_{2}$ with unique correct answers $y_{1}$ and $y_{2}$. Consider two models. The first model assigns probabilities $\hat{p}_{1}(y_{1}|x_{1})=1$ and $\hat{p}_{1}(y_{2}|x_{2})=0$. If we think of the probability a model assigns to an answer as its confidence in that answer, this model is highly confident and correct for problem $x_{1}$, but highly confident and wrong for $x_{2}$. Its pass@1 coverage is thus 50%. Moreover, since it always gets $x_{1}$ right and $x_{2}$ wrong, its pass@N coverage remains at 50%. In contrast, consider a second model which assigns probabilities $\hat{p}_{2}(y_{1}|x_{1})=\hat{p}_{2}(y_{2}|x_{2})=0.1$. This model has high confidence (0.9) but unfortunately on incorrect answers. Therefore its pass@1 accuracy is only 10%. However, it is willing to explore or hedge by placing some low confidence ($0.1$) on the other answer, which happens to be correct. Thus if this model samples $N$ times, the pass@N coverage increases with $N$, eventually approaching 100% as $N\rightarrow\infty$. Thus the first model outperforms the second in terms of pass@1 but not pass@N for large $N$. This indicates a tradeoff amongst policies: those that do better on pass@1 may not do better at pass@N. Of course the best one can do is be confident and correct, corresponding to the optimal model with $\hat{p}^{*}(y_{1}|x_{1})=\hat{p}^{*}(y_{2}|x_{2})=1$. However, this toy example reveals that if one cannot guarantee correctness, it can be beneficial to limit confidence and explore more solutions.
To demonstrate more generally the existence of tradeoffs between confident exploitation of a few answers versus unconfident exploration of many answers, as a function of the number of passes $N$, we prove two lemmas. To set up these lemmas, given any problem, let $\hat{p}_{i}$ denote the model probability or confidence assigned to answer $i$, and assume answers are sorted from highest to lowest confidence so that $\hat{p}_{i}\geq\hat{p}_{i+1},\forall\,i\geq 1$. Thus $\hat{p}_{1}$ is the model’s maximal confidence across all answers. Moreover, let $p_{i}$ be the probability across all the problems that the $i^{\text{th}}$ ranked answer is actually correct. Assume the model policy is approximately well calibrated so that higher confidence implies higher or equal probability of being correct, i.e. $p_{i}\geq p_{i+1},\forall\,i\geq 1$. We empirically verify that the model trained with CE loss on the MATH dataset approximately satisfies this assumption (Figure 12). Indeed, optimal policies maximizing pass@N coverage in Equation 2 are approximately well calibrated (See Lemma A.1). $p_{1}$ is then the model’s maximal accuracy across all answers. We prove:
@@
Assume a max accuracy $p_{1}$, and assume $\sum_{i=1}^{k}p_{i}\geq 1-\epsilon$ for some $10<\epsilon<$ 1. Then optimal policies maximizing pass@N coverage in Equation 2, subject to above accuracy and calibration constraints, must have max confidence upper bounded as
$\hat{p}_{1}^{*}\leq 1-\frac{k-1}{(k-1)^{\frac{N}{N-1}}p_{1}^{\frac{1}{N-1}}\left(1-p_{1}-\epsilon\right){}^{\frac{1}{1-N}}+1}.$
This upper bound is monotonically decreasing in $N$ (Figure 11), implying that at large $N$, an optimal policy for pass@N must limit its max confidence to be low, thereby favoring exploration and discouraging exploitation. Furthermore, we prove:
@@
Assume top two accuracies $p_{1}>p_{2}$. Then optimal policies maximizing pass@N coverage in Equation 2, subject to above accuracy and calibration constraints, must have max confidence obeying
@@
This lower bound is a monotonically decreasing function of $N$ (Figure 11), implying that at small $N$, optimal policies must have high max confidence, thereby favoring exploitation and discouraging exploration. Note in the limit $N\to 1^{+}$, the lower bound is always $1$, recovering the intuitive result that the optimal policy for pass@1 is to place 100% confidence on the highest accuracy answer.
@@ 4.3 Overconfidence prevents performance gains from scaling test-time compute
Figure 1: A model trained with CE loss becomes overconfident in its greedy completions, which harms its pass@N coverage; our proposed DCO objective limits this overconfidence. We fine-tune a Llama-3-8B base model on the MATH dataset to produce direct answers. $\hat{y}_{\text{greedy}}$ is the model’s greedy completion when choosing the most likely token at each sampling step. (a) The model trained with CE loss assigns progressively larger confidences $\hat{p}(\hat{y}_{\text{greedy}}|x)$ to its greedy completions over the course of training. (b) At the end of training, only a small portion of the model’s highly confident completions are correct. This will harm the model’s pass@N performance when scaling up $N$. (c) Same as (a) but shown for the DCO loss with $N^{\prime}=256$. The model trained on DCO shows a much milder overconfidence effect. (d) The confidence distribution of greedy completions after four epochs with DCO for various choices of $N^{\prime}$. As $N^{\prime}$ increases, the model’s confidence in its greedy completion is more stringently limited, directly as a consequence of the overconfidence regularizer $F$.
Lemma 4.2 suggests that at large $N$ it is beneficial to unconfidently explore by assigning low model confidence to many answers, while Lemma 4.3 suggests that at small $N$ it is beneficial to confidently exploit by assigning high model confidence to one or a few answers. These lemmas make a prediction that could explain the empirical observation in Table 1: namely, maximization of pass@1 makes the model overconfident, thereby preventing good performance on pass@N.
To test this theoretical prediction of model overconfidence, we estimated the max confidence of the model as $\hat{p}(y_{\mathrm{greedy}}|x)$ where $y_{\mathrm{greedy}}$ is the greedy completion to $x$ obtained by sampling autoregressively from the model $\hat{p}$, and at each step selecting the token with the highest probability. $\hat{p}(y_{\mathrm{greedy}}|x)$ approximates the max confidence $\hat{p}_{1}$ in Lemmas 4.2 and 4.3. For the model fine-tuned on MATH with CE loss, we plot the distribution of $\hat{p}(y_{\mathrm{greedy}}|x)$ over the test set in Figure 1 (a). This demonstrates the model becomes progressively more confident about its greedy completion over training, thereby confirming our theoretical prediction. In Figure 1 (b) we see why this is a problem: only some of the model’s highly confident answers are correct. Thus the model becomes overconfident and largely wrong. Scaling test-time compute cannot easily rescue such a model, as over multiple samples, it is likely to confidently provide the same wrong answers, explaining the origins of poor pass@N test performance. In Section C.4, we observe a similar trend of overconfidence on the out-of-distribution test set AIME24, further highlighting the generality of this failure mode.
@@ 5.1 A solution that naturally prevents overconfidence
is the $-\log$ probability that the model produces $y$ at least once in $N$ samples given $x$. Thus for each prompt $x$ and correct completion $y$ in the training set, we maximize the probability $y$ is found in $N$ passes. This loss naturally prevents model overconfidence, as can be seen via its gradient
@@
where $\ell_{\mathrm{CE}}$ is the standard CE loss on a single example, and $F(N,\hat{p}(y|x))=\frac{N(1-\hat{p}(y|x))^{N-1}\hat{p}(y|x)}{1-(1-\hat{p}(y|x))^{N}}$ is an extra overconfidence regularization factor that multiplies the standard CE gradient. Note that $F(N,\hat{p}(y|x))=1$ for $N=1$, so DCO reduces to CE for pass@1. Furthermore $F(N,\hat{p}(y|x))$ monotonically decreases in the model confidence $\hat{p}(y|x)$ (see Figure 2 (b)). Thus gradients for examples $(x,y)$ on which the model is more confident are attenuated, and this attenuation is stronger for larger $N$. This justifies the interpretation of $F(N,\hat{p}(y|x))$ as a regularizer that prevents model overconfidence. Indeed, for large $N$, $F(N,\hat{p}(y|x))\approx 0$ for confidence $\hat{p}(y|x)\gtrsim 1/N$. As soon as model confidence on an example exceeds $1/N$, its gradient becomes negligible. Thus interestingly, aligning training and test better through DCO naturally yields a simple emergent regularization of model overconfidence, which was itself identified as an impediment to performance gains through scaling test-time compute in Figure 1 (a). Indeed, $F(N,\hat{p}(y|x))$ is smaller for easier examples (Figure 13, right), effectively regularizing their contribution. This regularization mitigates the potential detrimental effects of overconfidence from easy examples as discussed in Section C.1.
@@
Figure 2: (a) DCO improves on CE loss for pass@N test coverage over a broad range of N and traces a Pareto-optimal frontier. We fine-tune Llama-3-8B base models on MATH to produce direct answers: one with CE loss and others using $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ for various $N^{\prime}$ (color-coded). Each curve shows pass@N coverage for a single fine-tuned model. Note that no $N^{\prime}$ is optimal for all $N$. The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. (See Figure 6, which highlights $N^{\prime}=256$, and  Figures 8 and 10 for Llama-3-70B and AIME24 results.) (b) DCO limits overconfidence by attenuating gradients for examples on which the model is highly confident. We plot the confidence regularization factor $F$ in Equation 4. For CE loss ($N=1$), $F=1$ regardless of confidence $\hat{p}(y|x)$. For $N>1$, $F$ decreases with $\hat{p}(y|x)$ and drops to zero around $1/N$—once confidence on an example reaches $O(1/N)$, $F$ vanishes, preventing it from further increasing. (c) Inverse confidence $N_{\text{eff}}$ at training time controls search-tree exploration shapes at test-time. We provide schematic proof search trees for models trained with DCOstep for small (left) and large (right) $N_{\mathrm{eff}}$. Circles are proof states; edges are tactics; solid edges are explored, while dashed edges are not. Small $N_{\mathrm{eff}}$ yields a narrow search tree that may miss correct proofs (red, dashed), while larger $N_{\mathrm{eff}}$ expands the tree to include the correct proof. However, if the tree becomes too wide ($N_{\mathrm{eff}}^{k}\gg N$ for a $k$-step proof), sampling the correct path becomes unlikely and pass@N coverage decreases as $\mathcal{C}^{N}\approx N/N_{\mathrm{eff}}^{k}$. Thus $N_{\mathrm{eff}}$, chosen at training time, is a powerful knob to control the tradeoff between exploitation and exploration at test time.

### v4

@@ 1 Introduction
  • We identify a misalignment between standard fine-tuning with CE loss and the pass@N coverage metric at test time (Section˜4.1).
  • We develop and experimentally verify a framework that suggests this misalignment arises from overconfidence induced by training on CE loss (Sections˜4.2 and 4.3).
  • We propose a loss function that directly optimizes the pass@N coverage metric, demonstrating consistent improvement over the CE loss objective and achieving superior accuracy frontiers on MATH and MiniF2F (Sections˜5.1, 5.2 and 5.3).
  • We extend our algorithm to more complex test-time scenarios including searching over proof-trees of varying shapes to improve automated theorem proving, and answering math questions using Chain-of-Thought reasoning traces, demonstrating improved mathematical reasoning in both cases (Sections˜5.3 and 5.4).
@@ 2 Related Works
Test-time compute and pass@N strategy. A growing body of work explores how training and test-time strategies interact to shape model performance. Jones [15] highlight a tradeoff between train and test-time compute in a controlled board game setting. Brown et al. [16] identify an exponentiated power law between coverage and number of in-context samples, captured phenomenologically in Levi [17] and Schaeffer et al. [18], while Snell et al. [1] explore compute-optimal strategies for scaling test-time inference. Our paper focuses primarily on the pass@N test-time strategy. Gui et al. [19] show that, in the best-of-N setting, the sample distribution is well-aligned with the reward model, and alignment methods have been proposed to distill this distribution and reduce sampling costs [19, 20]. However, their work focuses on improving the pass@1 performance from the Best-of-N policy, while we focus on directly improving the pass@N performance. Closely related to our work, Chow et al. [14] derives a similar training objective for reinforcement learning (RL) to directly optimize for the best-of-N test-time strategy. Li et al. [21] argue that fine-tuning with cross-entropy loss limits output diversity and propose a maximum entropy-based method to address this issue. Our work similarly identifies limitations of cross-entropy when scaling test-time compute, but focuses on the alignment of training and test-time strategies. In addition, Yue et al. [22] argue that fine-tuning with Group Relative Policy Optimization (GRPO) does not improve reasoning capabilities beyond the base model when evaluated with pass@N strategy. We will show that GRPO training may also lead to overconfidence in Figure˜3.
@@ 3 Problem setup
Minimizing the CE loss in Equation˜1 is equivalent to maximizing the pass@$1$ metric $\mathcal{C}_{\mathcal{D}}^{1}$ on a training set $\mathcal{D}$. But if we scale up test-time compute so that the pass@N metric $\mathcal{C}_{\mathcal{D}}^{N}$ on a test set $\mathcal{D}$ for $N\gg 1$ is the relevant performance metric, is $\mathcal{L}_{\mathrm{CE}}$ still a good training loss, or can we do better?
@@ 4.1 The CE loss induces overfitting for pass@N
To understand the impact of training with CE loss on pass@N test performance, we fine-tune Llama-3-8B-base [42] on the MATH [43] dataset. We start from the base model rather than Llama-3-8B-Instruct to avoid potential leakage of the MATH dataset into Llama-3-8B-Instruct through post-training. We follow Lightman et al. [24] and use $12,000$ problems for training and the remaining $500$ for testing. Here we train the model to provide a direct answer without a reasoning trace. We will discuss training with CoT in Section˜5.4. We confirm that the results presented in this section also hold for a larger model, Llama-3-70B-base, and direct readers to Section˜C.3 for additional results.
@@
Table˜1 reveals that the pass@N performance $\mathcal{C}_{\mathcal{D}}^{N}$ on a test set monotonically increases with the number of training epochs only for $N=1$, when minimizing CE loss is equivalent to maximizing $\mathcal{C}_{\mathcal{D}}^{1}$. However, for $N\geq 16$, minimizing CE loss during training does not monotonically increase $\mathcal{C}_{\mathcal{D}}^{N}$ at test; indeed for $N\geq 256$, pass@N test performance, remarkably, monotonically decreases with the number of training epochs, despite the fact that pass@1 performance monotonically increases. This effectively corresponds to a novel type of overfitting, in which test performance degrades over training, likely due to a mismatch between the test time (pass@N) and training time (pass@1) strategies.
@@ 4.2 Overfitting, confidence, and explore-exploit tradeoff
However, Lemma˜4.1 can fail when there is more than one problem. To understand this in a simple setting, consider a test set with two problems $x_{1}$ and $x_{2}$ with unique correct answers $y_{1}$ and $y_{2}$. Consider two models. The first model assigns probabilities $\hat{p}_{1}(y_{1}|x_{1})=1$ and $\hat{p}_{1}(y_{2}|x_{2})=0$. If we think of the probability a model assigns to an answer as its confidence in that answer, this model is highly confident and correct for problem $x_{1}$, but highly confident and wrong for $x_{2}$. Its pass@1 coverage is thus 50%. Moreover, since it always gets $x_{1}$ right and $x_{2}$ wrong, its pass@N coverage remains at 50%. In contrast, consider a second model which assigns probabilities $\hat{p}_{2}(y_{1}|x_{1})=\hat{p}_{2}(y_{2}|x_{2})=0.1$. This model has high confidence (0.9) but unfortunately on incorrect answers. Therefore its pass@1 accuracy is only 10%. However, it is willing to explore or hedge by placing some low confidence ($0.1$) on the other answer, which happens to be correct. Thus if this model samples $N$ times, the pass@N coverage increases with $N$, eventually approaching 100% as $N\rightarrow\infty$. Thus the first model outperforms the second in terms of pass@1 but not pass@N for large $N$. This indicates a tradeoff amongst policies: those that do better on pass@1 may not do better at pass@N. Of course the best one can do is be confident and correct, corresponding to the optimal model with $\hat{p}^{*}(y_{1}|x_{1})=\hat{p}^{*}(y_{2}|x_{2})=1$. However, this toy example reveals that if one cannot guarantee correctness, it can be beneficial to limit confidence and explore more solutions.
To demonstrate more generally the existence of tradeoffs between confident exploitation of a few answers versus unconfident exploration of many answers, as a function of the number of passes $N$, we prove two lemmas. To set up these lemmas, given any problem, let $\hat{p}_{i}$ denote the model probability or confidence assigned to answer $i$, and assume answers are sorted from highest to lowest confidence so that $\hat{p}_{i}\geq\hat{p}_{i+1},\forall\,i\geq 1$. Thus $\hat{p}_{1}$ is the model’s maximal confidence across all answers. Moreover, let $p_{i}$ be the probability across all the problems that the $i^{\text{th}}$ ranked answer is actually correct. Assume the model policy is approximately well calibrated so that higher confidence implies higher or equal probability of being correct, i.e. $p_{i}\geq p_{i+1},\forall\,i\geq 1$. We empirically verify that the model trained with CE loss on the MATH dataset approximately satisfies this assumption (Figure˜13). Indeed, optimal policies maximizing pass@N coverage in Equation˜2 are approximately well calibrated (See Lemma˜A.1). $p_{1}$ is then the model’s maximal accuracy across all answers. We prove:
@@
Assume a max accuracy $p_{1}$, and assume $\sum_{i=1}^{k}p_{i}\geq 1-\epsilon$ for some $0<\epsilon<1-p_{1}$. Then optimal policies maximizing pass@N coverage in Equation˜2, subject to the above accuracy and calibration constraints, have a confidence upper bound $u(N,p_{1},\epsilon,k)$, which is a non-increasing function of $N$.
We proved this upper bound as well as the proof in Appendix˜A. This upper bound is monotonically decreasing in $N$ (Figure˜12), implying that at large $N$, an optimal policy for pass@N must limit its max confidence to be low, thereby favoring exploration and discouraging exploitation. Furthermore, we prove:
@@
Assume top two accuracies $p_{1}>p_{2}$. Then optimal policies maximizing pass@N coverage in Equation˜2, subject to above accuracy and calibration constraints, must have max confidence obeying
@@
This lower bound is a monotonically decreasing function of $N$ (Figure˜12), implying that at small $N$, optimal policies must have high max confidence, thereby favoring exploitation and discouraging exploration. Note in the limit $N\to 1^{+}$, the lower bound is always $1$, recovering the intuitive result that the optimal policy for pass@1 is to place 100% confidence on the highest accuracy answer.
@@ 4.3 Overconfidence prevents performance gains from scaling test-time compute
Figure 1: A model trained with CE loss becomes overconfident in its greedy completions, which harms its pass@N coverage; our proposed DCO objective limits this overconfidence. We fine-tune a Llama-3-8B base model on the MATH dataset to produce direct answers. $\hat{y}_{\text{greedy}}$ is the model’s greedy completion when choosing the most likely token at each sampling step. (a) The model trained with CE loss assigns progressively larger confidences $\hat{p}(\hat{y}_{\text{greedy}}|x)$ to its greedy completions over the course of training. (b) At the end of training, only a small portion of the model’s highly confident completions are correct. This will harm the model’s pass@N performance when scaling up NN. (c) Same as (a) but shown for the DCO loss with $N^{\prime}=256$. The model trained on DCO shows a much milder overconfidence effect. (d) The confidence distribution of greedy completions after four epochs with DCO for various choices of $N^{\prime}$. As $N^{\prime}$ increases, the model’s confidence in its greedy completion is more stringently limited, directly as a consequence of the overconfidence regularizer $F$.
Lemma˜4.2 suggests that at large $N$ it is beneficial to unconfidently explore by assigning low model confidence to many answers, while Lemma˜4.3 suggests that at small $N$ it is beneficial to confidently exploit by assigning high model confidence to one or a few answers. These lemmas make a prediction that could explain the empirical observation in Table˜1: namely, maximization of pass@1 makes the model overconfident, thereby preventing good performance on pass@N.
To test this theoretical prediction of model overconfidence, we estimated the max confidence of the model as $\hat{p}(y_{\mathrm{greedy}}|x)$ where $y_{\mathrm{greedy}}$ is the greedy completion to $x$ obtained by sampling autoregressively from the model $\hat{p}$, and at each step selecting the token with the highest probability. $\hat{p}(y_{\mathrm{greedy}}|x)$ approximates the max confidence $\hat{p}_{1}$ in Lemmas˜4.2 and 4.3. For the model fine-tuned on MATH with CE loss, we plot the distribution of $\hat{p}(y_{\mathrm{greedy}}|x)$ over the test set in Figure˜1 (a). This demonstrates the model becomes progressively more confident about its greedy completion over training, thereby confirming our theoretical prediction. In Figure˜1 (b) we see why this is a problem: only some of the model’s highly confident answers are correct. Thus the model becomes overconfident and largely wrong. Scaling test-time compute cannot easily rescue such a model, as over multiple samples, it is likely to confidently provide the same wrong answers, explaining the origins of poor pass@N test performance. In Section˜C.4, we observe a similar trend of overconfidence on the out-of-distribution test set AIME24, further highlighting the generality of this failure mode.
@@ 5.1 A solution that naturally prevents overconfidence
is the $\log$ probability that the model produces $y$ at least once in $N$ samples given $x$. Thus for each prompt $x$ and correct completion $y$ in the training set, we maximize the probability $y$ is found in $N$ passes. This loss naturally prevents model overconfidence, as can be seen via its gradient
@@
where $\ell_{\mathrm{CE}}$ is the standard CE loss on a single example, and $F(N,\hat{p}(y|x))=\frac{N(1-\hat{p}(y|x))^{N-1}\hat{p}(y|x)}{1-(1-\hat{p}(y|x))^{N}}$ is an extra overconfidence regularization factor that multiplies the standard CE gradient. Note that $F(N,\hat{p}(y|x))=1$ for $N=1$, so DCO reduces to CE for pass@1. Furthermore $F(N,\hat{p}(y|x))$ monotonically decreases in the model confidence $\hat{p}(y|x)$ (see Figure˜2 (b)). Thus gradients for examples $(x,y)$ on which the model is more confident are attenuated, and this attenuation is stronger for larger $N$. This justifies the interpretation of $F(N,\hat{p}(y|x))$ as a regularizer that prevents model overconfidence. Indeed, for large $N$, $F(N,\hat{p}(y|x))\approx 0$ for confidence $\hat{p}(y|x)\gtrsim 1/N$. As soon as model confidence on an example exceeds $1/N$, its gradient becomes negligible. Thus interestingly, aligning training and test better through DCO naturally yields a simple emergent regularization of model overconfidence, which was itself identified as an impediment to performance gains through scaling test-time compute in Figure˜1 (a). Indeed, $F(N,\hat{p}(y|x))$ is smaller for easier examples (Figure˜14, right), effectively regularizing their contribution. This regularization mitigates the potential detrimental effects of overconfidence from easy examples as discussed in Section˜C.1.
@@
Figure 2: (a) DCO improves on CE loss for pass@N test coverage over a broad range of N and traces a Pareto-optimal frontier. We fine-tune Llama-3-8B base models on MATH to produce direct answers: one with CE loss and others using $\mathcal{L}_{\text{DCO}}^{N^{\prime}}$ for various $N^{\prime}$ (color-coded). Each curve shows pass@N coverage for a single fine-tuned model. Note that no $N^{\prime}$ is optimal for all $N$. The black curve is a Pareto-optimal performance frontier traced by the max of coverage curves for DCO over all $N^{\prime}$. (See Figure˜6, which highlights $N^{\prime}=256$, and  Figures˜8 and 10 for Llama-3-70B and AIME24 results.) (b) DCO limits overconfidence by attenuating gradients for examples on which the model is highly confident. We plot the confidence regularization factor $F$ in Equation˜4. For CE loss ($N=1$), $F=1$ regardless of confidence $\hat{p}(y|x)$. For $N>1$, $F$ decreases with $\hat{p}(y|x)$ and drops to zero around 1/N1/N—once confidence on an example reaches $O(1/N)$, $F$ vanishes, preventing it from further increasing. (c) Inverse confidence $N_{\text{eff}}$ at training time controls search-tree exploration shapes at test-time. We provide schematic proof search trees for models trained with DCOstep for small (left) and large (right) $N_{\mathrm{eff}}$. Circles are proof states; edges are tactics; solid edges are explored, while dashed edges are not. Small $N_{\mathrm{eff}}$ yields a narrow search tree that may miss correct proofs (red, dashed), while larger $N_{\mathrm{eff}}$ expands the tree to include the correct proof. However, if the tree becomes too wide ($N_{\mathrm{eff}}^{k}\gg N$ for a $k$-step proof), sampling the correct path becomes unlikely and pass@N coverage decreases as $\mathcal{C}^{N}\approx N/N_{\mathrm{eff}}^{k}$. Thus $N_{\mathrm{eff}}$, chosen at training time, is a powerful knob to control the tradeoff between exploitation and exploration at test time.